OpenStudy (anonymous):

f(x)=4x^2+3x+4. How do I solve this for -f(x)?

jimthompson5910 (jim_thompson5910):

-f(x) is just the negation of everything so if f(x) = x+2 for instance, then f(x) = x+2 -f(x) = -(x+2) ... Multiply both sides by -1. -f(x) = -x - 2 ... distribute

OpenStudy (anonymous):

Thanks :)

jimthompson5910 (jim_thompson5910):

yw